OpenStudy (anonymous):

Find the first four iterates of the function f(z)=z^2+2+i with the initial value of z0=2+2i

OpenStudy (anonymous):

If this is supposed to be a sequence, with z representing the previous term, what happens when you apply the function to your initial value?

OpenStudy (anonymous):

I appreciate your trying to help . . . but, my way of learning doesn't involve being asked questions in order to get it out of me.

OpenStudy (anonymous):

I'm just trying to figure out what's going on here, actually. Is this supposed to be a sequence? (honest question)

OpenStudy (anonymous):

Sorry. I honestly, have no clue. I'm so confused with it myself :/

OpenStudy (anonymous):

I think it is, so I'll rewrite it as if it were to make it clear to myself. Then, we can both pretend I'm right. \[f(z_n)=z _{n-1}^{2}+2-i\]

OpenStudy (anonymous):

Do you know how to square complex numbers?

OpenStudy (anonymous):

Lol, alright. & nope, I don't.

OpenStudy (nurali):

f( z) = z^2 + 2 + i , z0 = 2+2i Remember the formulas: (a ± b)^2 = a^2 ± 2ab + b^2 Now what if we replace "b" with "bi" then: (a ± bi)^2 = a^2 ± 2abi + b^2 i^2 >> but i^2 = - 1 by definition so: (a ± bi)^2 = a^2 ± 2abi - b2 ----> This is the key to solve this problem z1 = f(z0) = (2 + 2i)^2 + 2 + i = 4 + 8i +(4i^2) + 2 + i = 4 - 4 + 2 + 9i = 2 + 9i z2 = f(z1) = (2 + 9i)^2 + 2 + i = 4 + 36i - 81 + 2 + i = -75 + 37i z3 = f(z2) = (-75 + 37i)^2 + 2 + i = 5625 - 5550i - 1369 + 2 + i = 4258 - 5549i z4= f(z3) = (4258 - 5549i)^2 + 2 + i = 18130564 - 47255284i - 30791401 + 2 + i = -12660835 - 47255283i

OpenStudy (anonymous):

Thank you :) Both, of you.

OpenStudy (nurali):

Anytime.